these are the coordinate of a triangle determine wether it is scalene, isosceles or equilateral. T(1,1) R(3,8) S(6,4)
Answer by stanbon(75887)

these are the coordinate of a triangle determine wether it is scalene, isosceles or equilateral. T(1,1) R(3,8) S(6,4)
TR = sqrt[(3-1)^2+(8-1)^2] = sqrt[4 + 49] = sqrt(53)
------------
TS = sqrt[5^2 + 3^2] = sqrt[34]
-------------
RS = sqrt[3^2 + 4^2] = 5
-----------------------------
Ans: scalene
-------------
Cheers,
Stan H.
------------